# Quantum Error Correction resilient against Atom Loss

### v1

@@ 8 Conclusion
In conclusion, this study demonstrates that the integration of LDUs with the surface code enables efficient QEC in the presence of both atom loss and depolarizing noise, provided losses and depolarizing errors remain below the threshold. Specifically, for a surface code of distance $d$ and no depolarizing noise, it is possible to correct up to $d-1$ errors, similar to an erasure channel, even when only partial knowledge of the loss location is available through the LDUs. This makes the system highly resilient to atom loss.
@@
Several research directions could be explored to further refine and improve the error models and the error correction protocol for neutral atom quantum processors. One direction is to investigate more realistic noise models by including imperfect measurements, noisy single-qubit gates or customized CZ gate Pauli noise channels tailored specifically for neutral atoms, providing a more accurate representation of the noise environment in these systems [56]. Additionally, recent studies have shown that leakage out of the computational subspace can be converted to erasure channel [22, 26, 21], which enhance the overall error correction capability of the system.
Another direction is to complexify the loss model by distinguishing losses due to heating and collisions, typically occurring between gate applications, from those during CZ gates caused by anti-trapping effects. The former type of loss should be proportional to the time spent between two consecutive CZ gates while the latter may impact the remaining atom involved in the CZ gate; however due to the Rydberg blockade effect, it is unlikely that both atoms would be simultaneously lost. Moreover, an imperfect reloading scheme could be considered, where atom loss might occur during shuttling.
@@
The present analysis can be readily extended to investigate different QEC protocols for neutral atom quantum computing such as the recently introduced LDPC codes [51, 50, 57, 58]
@@ Acknowledgements
S. J. gratefully acknowledges discussions with Mikhail Lukin. This work was founded by the European Union’s Horizon Europe program HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa) and by the Horizon 2020 program under the Marie Sklodowska-Curie project 955479 (MOQS) and by the French National Research Agency under the Investments of the Future Program project ANR-21-ESRE-0032 (aQCess).
@@ Appendix A Conditional probabilities used in the “loss-aware” decoder
In this Appendix, we explicitly compute the conditional probabilities of the “loss-aware” decoder presented in Section 4 based on Eq. (2).
@@ Appendix E Logical error for the naive decoder
$\sum_{k=\lceil\frac{d+1}{4}\rceil}^{\frac{d+1}{2}}a_{k}(d)p_{l}^{k}$ (35)
where $a_{k}(d)$ are fitting parameters which depends on the distance code. This functional form allows us to accurately capture the behavior of the logical error probability when using the naive decoder.

### v2

@@ 9 Conclusion
In conclusion, this study demonstrates that the integration of LDUs with the surface code enables efficient QEC in the presence of both atom loss and depolarizing noise, provided losses and depolarizing errors remain below the threshold. Specifically, for a surface code with distance $d$ and no depolarizing noise, it is possible to correct up to $d-1$ errors under the first loss model, where the loss of an atom does not affect the neighboring qubit. This behavior is similar to an erasure channel, although only partial knowledge of the loss location is available through the LDUs. When the loss impacts neighboring atoms, the scaling law is slightly affected, with the logical error probability scaling approximately with a power law of exponent $\sim d-2$. For both loss models, the system is highly resilient to atom loss.
@@
Several research directions could be explored to further refine and improve the error models and the error correction protocol for neutral atom quantum processors. One direction is to investigate more realistic noise models by including imperfect measurements, noisy single-qubit gates or customized CZ gate Pauli noise channels tailored specifically for neutral atoms, providing a more accurate representation of the noise environment in these systems [66]. Additionally, recent studies have shown that leakage out of the computational subspace can be converted to erasure channel [22, 27, 21, 29], which enhance the overall error correction capability of the system.
Another direction is to complexify the loss model by distinguishing losses due to heating and collisions, typically occurring between gate applications, from those during CZ gates caused by anti-trapping effects. The former type of loss should be proportional to the time spent between two consecutive CZ gates while the latter may impact the remaining atom involved in the CZ gate by enhancing leakage mechanism. Two directions can be explored: (i) assume the ability to repump the leaked atom back into the computational state, such as considered in Sec. 8 where the remaining atom experienced a maximally biased Z error after RC or (ii) consider leakage-to-loss conversion schemes which may introduce correlated losses between atoms involved in the same CZ gate (see Supplementary Information of [27]). Moreover, an imperfect reloading scheme could be considered, where atom loss might occur during shuttling.
@@
The present analysis can be readily extended to investigate different QEC protocols for neutral atom quantum computing such as the recently introduced LDPC codes [60, 67, 68, 69, 61, 70]
@@ Acknowledgements
S. J. gratefully acknowledges discussions with Mikhail Lukin. H. P. and G. P. would like to express their gratitude for insightful discussions with Jeff Thompson. This work was founded by the European Union’s Horizon Europe program HORIZON-CL4-2021-DIGITAL-EMERGING-01-30 via the project 101070144 (EuRyQa) and by the Horizon 2020 program under the Marie Sklodowska-Curie project 955479 (MOQS) and by the French National Research Agency under the Investments of the Future Program project ANR-21-ESRE-0032 (aQCess).
Data and code availability
Data and the code used to generate detector error models, simulate quantum circuits, compute the logical error rates and analyze the results are available on Zenodo open repository
@@ Appendix A Conditional probabilities used in the “loss-aware” decoder
In this Appendix, we explicitly compute the conditional probabilities of the “loss-aware” decoder presented in Sec. 4 based on Eq. (4).
@@ Appendix E Logical error for the naive decoder
$\sum_{k=\lceil\frac{d+1}{4}\rceil}^{\frac{d+1}{2}}b_{k}(d)p_{l}^{k}$ (37)
where $b_{k}(d)$ are fitting parameters which depends on the distance code. This functional form allows us to accurately capture the behavior of the logical error probability when using the naive decoder.